### v1

# Renovating Names in Open-Vocabulary Segmentation Benchmarks

Abstract
Names are essential to both human cognition and vision-language models. Open-vocabulary models utilize class names as text prompts to generalize to categories unseen during training. However, name qualities are often overlooked and lack sufficient precision in existing datasets. In this paper, we address this underexplored problem by presenting a framework for “renovating” names in open-vocabulary segmentation benchmarks (RENOVATE). Through human study, we demonstrate that the names generated by our model are more precise descriptions of the visual segments and hence enhance the quality of existing datasets by means of simple renaming. We further demonstrate that using our renovated names enables training of stronger open-vocabulary segmentation models. Using open-vocabulary segmentation for name quality evaluation, we show that our renovated names lead to up to 16% relative improvement from the original names on various benchmarks across various state-of-the-art models. We provide our code and relabelings for several popular segmentation datasets (ADE20K, Cityscapes, PASCAL Context) to the research community.
Keywords:
Vision-Language Datasets Open-Vocabulary Segmentation Renaming
Categorizations which humans make of the concrete world are not arbitrary but highly determined. In taxonomies of concrete objects, there is one level of abstraction at which the most basic category cuts are made.
— Eleanor Rosch, Basic Objects in Natural Categories [36]
1 Introduction
We use names every day. Imagine wandering the trails of a national park – do you pause by a “lake” or simply a body of “water”? While driving through urban sprawl, do you see “trees” or “vegetation” along the road? Our instinctual use of terms like “lake” and “trees” illustrates the human propensity to categorize the world around us into “basic categories” — rich in information and readily identifiable without unnecessary complexity [37, 36, 40]. Such categorization is fundamental to human cognition and communication.
In stark contrast, recent advancements in open-vocabulary models [14, 11, 9, 49, 20, 12, 16, 54, 17] struggle to replicate this nuanced aspect of human categorization. While these models have made strides in generalizing to both familiar and novel categories via textual prompts, they are often hampered by the imprecise and sometimes wrong names provided in existing benchmarks. In fact, most datasets are labeled with class names that serve merely as identifiers to distinguish classes within a dataset, rather than descriptive labels aligning with the “basic categories” that match with the actual visual contents. As shown in Fig. 1, existing names are often inaccurate, too general, or lack enough context, leading to discrepancies between the model’s outputs and the actual visual segments. This misalignment indicates a pressing need for a reassessment and refinement of name labeling practices. By adopting a naming scheme that aligns more closely with human categorization, we can pave the way for enhanced open-vocabulary generalization and more accurate model evaluation.
(a) Original names can be inaccurate.
(b) Original names can be too general.
(c) Original names can lack contexts.
Figure 1: Problems of names in current segmentation benchmarks. We demonstrate with examples from well-known datasets: MS COCO [22], ADE20K [57], and Cityscapes [7]. Our renovated names are more aligned with the visuals and help models to generalize better.
In this work, we focus on renovating the names for open-vocabulary segmentation benchmarks, as their dense annotations pose a greater challenge and offer broader applicability than other recognition tasks. The importance of names in open-vocabulary segmentation is often overlooked, with one exception, OpenSeg [11], which manually inspected and modified the class names of several segmentation benchmarks, resulting in considerable performance gains of their proposed model. While their work demonstrated the importance of choosing good names, a manual approach is subjective and difficult to scale to finer granularity. To date, there exists no systematic study on how to rename benchmarks in a scalable and principled way. Our work is the first attempt to tackle this challenge.
In this work, we introduce a scalable, simple, yet general method for renaming segmentation benchmarks. Our approach leverages foundation models to automate the renaming process, significantly reducing the manual labor traditionally required. Towards this goal, our method can be divided into two steps. First, we narrow down the name search space from the whole language space to a curated list of candidate names. This list is generated by leveraging the original class names with contextually relevant nouns extracted from visual contents by an image captioning model, thereby streamlining the search for names. Next, we employ a specially trained renaming model to identify and select the best-matching candidate name for every ground-truth segmentation mask. In this way, we match an individual name for each instance without any extra human annotations.
To validate the quality of our renovated names, we first conduct a careful human preference study showing that our method successfully finds more accurate names for current segmentation benchmarks. We then apply our renovated names to upgrade existing benchmarks, resulting in datasets that not only contain more precise names of the visual segments, but also expand the number of classes by $4\sim 5$ times. This enhancement makes the benchmarks significantly more challenging and realistic than the original benchmarks. Then, we use our renovated names for training open-vocabulary models and show that our names improve the open-vocabulary generalization performances on various benchmarks. Finally, we leverage pre-trained open-vocabulary models to automate the name quality evaluation process. Our results on the ADE20K, Cityscapes, and PASCAL Context benchmarks show that the renovated names improve pre-trained models up to 16% over the original names, suggesting that our names match more accurately with the visual segments.
In summary, our work makes the following contributions:
  • We point out the name problems in existing open-vocabulary segmentation benchmarks.
  • We propose a scalable, simple, general framework (RENOVATE) for renaming segmentation benchmarks.
  • We demonstrate that our renovated names are of high quality and can upgrade existing benchmarks to be more challenging and realistic, as well as help train models with stronger open-vocabulary capabilities.
2 Related Work
Open-Vocabulary Segmentation Methods. Prior work in open-vocabulary segmentation [56, 2, 47, 50, 11, 9, 49, 20, 5] focused on adapting vision-language models (VLMs) [3, 41, 32] to the segmentation task without compromising pre-trained vision-langauge alignment. These efforts have concentrated mainly on the architecture design, exploring different backbones [48, 52], mask proposal branch designs [50, 11], and mask feature extraction techniques [49, 30, 42]. Our work differs from them in that we aim to adapt pre-trained vision-language representations for segment-wise name matching through our novel designs like candidate name generation techniques and our specialized renaming model.
Names in Open-Vocabulary Segmentation Benchmarks. Despite its apparent importance, the naming problem in open-vocabulary segmentation benchmarks has not received adequate attention. An exception is the work by [29], which proposes to learn class-specific word embeddings for improved model adaptation to new datasets. However, these learned embeddings are not in the language space and cannot be used by other models. To the best of our knowledge, OpenSeg [11] is the only work that proposes to improve the quality of the names. However, their approach relies on manual inspection of the datasets, which can be subjective and hard to scale. In this paper, we systematically explore the renaming challenge, culminating in a scalable, simple, yet general framework to overcome these limitations.
Original name Context names Candidate names field lush, field, sky, green, grass, tree, road, hillside, grassy, rural rural field, roadside field, green field, crop field, sports field, grassland, grassy hillside box room, stool, chair, floor, market, table, food, lamp, paper storage box, packing box, file box, box container decorative box, display box, paper box, food box cradle room, infant bed, nursery, dresser, carpet, bedroom, chair, bureau, lamp, armchair bedroom cradle, cradle, infant bed, nursery cradle, baby cradle, infant cradle, wooden cradle
Table 1: Examples of context names and generated candidate names for three selected classes from ADE20K. Context names are key to comprehending general terms such as “field” and “box” and disambiguating polysemous terms like “cradle”, which, in this context, refers to a baby bed rather than a phone cradle or a mining tool.
3 Method
Our method RENOVATE aims to rename the original names in a given dataset with ground-truth mask annotations. As shown in Fig. 2, we first generate a pool of candidate names, covering potential variants of the original class names. Then, we train a specially designed renaming model that can measure the alignment between segments in pixel space and names in language space. Equipped with the renaming model, we select the top-matching names from the candidate pool for each segment, thereby producing segment-level, refined names that offer enhanced accuracy and details in characterizing the dataset, see Fig. 3 and Fig. 4.
3.1 Generating candidate names
We use GPT-4 [28] for creating a pool of candidate names. A naive solution is to prompt GPT-4 with the original name and ask it to generate synonyms and hierarchical concepts. However, since the original names are often too general and ambiguous, GPT-4 does not have sufficient knowledge to generate high-quality candidates. Therefore, we propose to exploit the visual contents for generating some “context names” (Tab. 1) that assist GPT-4 in comprehending the category’s meaning prior to generating candidate names.
Figure 2: Overview of candidate name generation and renaming model training. We generate candidate names based on the context names and train the renaming model to match them with the segments. For illustration clarity, we show only one segment. In practice, multiple segments are jointly trained, pairing with the text queries.
As shown in Fig. 2, for each category, we use an image captioning method to process all training images that contain that specific category based on ground-truth annotations. From the generated captions, we further extract nouns by text parsing and filtering as done in CaSED [6]. We observe that nouns appearing more frequently tend to offer deeper insights into the typical environments or traits associated with the category. Therefore, we construct context names by selecting the top 10 most recurrent nouns for each category. We then use them as additional inputs alongside the original class names to prompt GPT-4.
Specifically, we instruct GPT-4 to focus on two types of candidate name generation: (1) Synonyms or subcategories, which can expand the lexical scope of the overly general class name. For instance, the class name “grass” could yield variations like “lawn” or “turf”. For this type, synonyms and subcategories are considered together, acknowledging their often interchangeable nature in certain contexts. (2) Providing short, distinct contexts to polysemous class names. As an example, the ambiguous class name “fan” could be elaborated into “ceiling fan” or “floor fan”, offering clearer specificity. As shown in Tab. 1, the 5-10 candidate names generated by GPT-4 are more precise than the original names, reflecting the context information encoded by the context names. We provide detailed GPT-4 prompts in the supplement.
3.2 Training for candidate name selection
When using GPT-4 to generate candidate names, generating a surplus is not an issue. Among the generated candidates, we will only keep those that are better aligned with the visual contents from the images. Particularly, we train a model that can assess the alignment between each name and segment. For training such a model, we use segmentation as the proxy task. Conceptually, a candidate name that well describes the segment in the image should help the segmentation model recover its ground-truth mask and make a correct classification. Therefore, the model should allow vision-language interaction, and the supervision signal should encourage to use textual information for mask prediction and classification.
Our renaming model is illustrated in Fig. 2. Its vision part starts from a CLIP-based vision backbone followed by a pixel decoder that gradually upscales the backbone features to generate per-pixel embeddings $F_{\text{pix}}\in\mathbb{R}^{H\times W\times C}$. The candidate names of the categories present in the image are processed by the CLIP text encoder. The interaction of textual and visual information takes place at the transformer decoder. The initial queries $X^{(0)}$ are from the text embeddings, and then updated by the transformer decoder consisting of $L$ blocks using the multi-scale visual features from the pixel decoder. At the output of each block, the queries $X^{(l)}$ generate intermediate mask and class predictions as follows:
$\hat{M}^{(l)}=\mathrm{sigmoid}(X^{(l)}*F_{\text{pix}}),\;\;\;\hat{c}^{(l)}\;=\mathrm{softmax}\left[\mathrm{Linear}(X^{(l)})\right]$ (1)
where *** refers to pixel-wise multiplication. The final predictions are made by the queries $X^{(L)}$. Our pixel decoder, mask prediction and classification follow the design of Mask2Former [4], which handles panoptic, instance, and semantic segmentation in a unified manner. The key differences lie in 1) the transformer decoder design, and 2) the training strategy, which we detail subsequently.
Transformer Decoder. Our transformer decoder uses text embeddings as input queries, unlike most open-vocabulary segmentation models that employ text embeddings solely for mask classification as weights of the Linear layer in Eq. 1. This key change fosters an earlier integration of text and visual features at varying scales, essential for leveraging text information in segmentation tasks. As shown in Fig. 2, the text embedding of each candidate name from the segment’s class “bicycle” feeds into the transformer decoder, generating corresponding mask and class predictions for each candidate name.
Next, within each block of the transformer decoder, the masked cross-attention (CA) layer adopts ground-truth masks as the attention biases, and the self-attention (SA) layer [45] and feed-forward network (FFN) are the same as in Mask2Former. The queries interact with the visual features at the masked CA layers, which attend within the region defined by the ground-truth mask, i.e., the attention bias. In the example of Fig. 2, the visual features corresponding to the segment “bicycle” can therefore contribute more effectively to refining the query. To encourage the model to rely more on the textual information in prediction, we further randomly replace the ground-truth masks with predicted masks from the preceding block. Importantly, while ground-truth mask based attention biases augment the V-L interaction by focusing on the precise segment regions, they do not reduce the task to a straightforward use of ground-truth mask inputs for mask prediction, as the output queries remain a set of vectors. The final mask prediction $\hat{M}^{(L)}$ is based on the correlation of these output queries $X^{(L)}$ with the feature map $F_{\text{pix}}$ generated by the pixel decoder as in Eq. 1.
Training Strategy. With both the vision and text encoder of CLIP kept frozen, we train the transformer decoder together with the pixel decoder for segmentation. To recover the ground-truth mask $M_{i}$ and class $c_{i}$, the transformer decoder makes multiple mask and class predictions with candidate names from the class $c_{i}$. The prediction with the highest Intersection over Union (IoU) with $M_{i}$ is selected for loss computation:
$\mathcal{L}_{i}=\mathcal{L}_{\text{mask}}(M_{i},\hat{M}_{\mathrm{best}})+\mathcal{L}_{\text{class}}(c_{i},\hat{c}_{\mathrm{best}}),$ (2)
where $\mathcal{L}_{\text{mask}}$ and $\mathcal{L}_{\text{class}}$ are the mask localization and classification loss functions in Mask2Former. To provide extra supervision on the name quality, we append the candidate names with a “negative” name randomly selected from a different category than the ground-truth class $c_{i}$. If a “negative” name scores the highest IoU, we supervise the predictions with an empty mask and the “void” class ($c_{void}$) which is one extra class in addition to the training classes:
$\mathcal{L}_{i}=\mathcal{L}_{\text{mask}}(\mathbf{0},\hat{M}_{\mathrm{best}})+\mathcal{L}_{\text{class}}(c_{void},\hat{c}_{\mathrm{best}}).$ (3)
The “void” class was also used in Mask2Former classification when the prediction mask is not matched to any ground-truth mask in the image. Having both the positive and negative supervision, we effectively incentivize the model to favor names of high quality and penalize those of low quality, thereby aiding in the accurate identification of the best-matching names for each segment.
(a) Process of obtaining renovated names.
(b) Renovated name distribution within the original “person” class.
Figure 3: Obtaining renovated names. In (a) we illustrate how we use the renaming model to obtain a renovated name for each segment. In (b) we demonstrate that the renaming results are helpful to dataset analysis with examples from “person” class.
Figure 4: Examples of renovated names on segments from the validation sets of ADE20K and Cityscapes. For each segment, we show the original name below the image and the renovated name in the text box. See more visual results in the supplements.
3.3 Obtaining renovated names
As illustrated in Fig. 2(a), to obtain the renovated names, we run the trained renaming model to associate each ground-truth mask with the candidate name. For each segment, our renaming model first ranks the names by the IoU scores between the predicted masks and the ground-truth mask, then selects the top-scoring name as the renovated name. While the IoU ranking scores are not probabilities, they are valuable to understand the relative and absolute model confidence in the names, i.e., how the chosen name stands against the other candidates and whether the selection is indeed confident. In the experiments (Sec. 4.2), we make use of this information to speed up human verification. We show some examples of renovated names in Fig. 4 and more in the supplements.
Our renovated names are readily usable for analyzing the dataset in a more fine-grained manner, as exemplified in Fig. 2(b). This can be helpful for countering the dataset biases during dataset curation and collection. Note that our renovated names in each original class are not necessarily mutually exclusive. In fact, we intentionally did not include the mutual exclusivity as a constraint during candidate name generation because different visual segments may be matched to names at different hierarchies which are not mutually exclusive. For example, in Fig. 4, a person is named as “engineer” (second row), likely due to the clothing, while the other person is named as “woman” (third row), possibly from the physical appearance. But “engineer” and “woman” are not mutually exclusive names. In the future, it would be an interesting extension to further study the hierarchies and other relationships of different names in a dataset.
4 Experiments
Datasets. We apply RENOVATE to MS COCO [22], ADE20K [57], Cityscapes [7], and PASCAL Context [27]. Both MS COCO and Cityscapes have panoptic annotations. ADE20K has two annotation sets: ADE-150 with panoptic annotations for 150 classes and ADE-847 with only semantic annotations for 847 classes. PASCAL Context has only semantic annotations for 459 classes. When the panoptic annotations are available, we perform RENOVATE on panoptic segments, otherwise, we rename the semantic segments.
Implementation details. To generate context names, our default image captioning model is CaSED [6], which retrieves captions by matching vision embeddings with text embeddings of captions from a large-scale external image-caption dataset PMD [39]. For each dataset to be renamed, we train a renaming model for 60k iterations with a batch size of 16 only on the training set. See details in the supplements.
Name baselines. We consider the only two other existing name baselines for open-vocabulary segmentation benchmarks: (1) Original class names; (2) OpenSeg class names, which were manually curated by previous researchers [11].
(a) User interface for human preference study.
(b) Human preference results.
Figure 5: Human preference study. A survey of 20 researchers is conducted to compare preferences between the original names versus RENOVATE names on the validation sets. RENOVATE names are favored in 82% cases.
4.1 Human preference study
As demonstrated in Fig. 5, we first study the name quality of our renovated names with a human preference study. Specifically, we first randomly select 10% of all the segments in the validation set of each dataset. Then for each segment, we ask 20 researchers unfamiliar with the project about their preference between its original name and our model-selected name. Our results demonstrate that our renovated names are preferred in 82% cases, showing a clear advantage over the original names. We provide more details of our human study in the supplements.
4.2 Upgrading benchmarks with renovated names
Setup. We demonstrate that renovated names can be used to upgrade existing benchmarks on two popular open-vocabulary benchmarks, ADE20K (ADE-150) and Cityscapes. Thanks to our renovated names, our upgraded benchmarks contain 4 to 5 times more classes than the original ones. Besides being more fine-grained, our benchmarks are also more realistic, aligning more closely with how humans use names to prompt models. As we show later, these renovated names provide valuable extra information to the existing benchmarks.
To ensure high quality names for segments in the validation sets, we involve five human annotators in a rigorous verification process. Specifically, we present human annotators with the top 3 name suggestions along with their confidence scores from our renaming model. We then ask them to select the best-matching names among them. For the few segments that do not match with any of the top 3 names, the annotators will check the whole candidate name list of the corresponding groundtruth class. In case of disagreements over the name selection, all the annotators discuss and decide the best choices. We will release these annotations upon publication.
Along with the benchmarks, we incorporate new evaluation metrics [58] that account for the semantic similarity between names, rather than the standard evaluation metrics of panoptic quality (PQ), Average Precision (AP), and mean Intersection-over-Union (mIoU) which penalize the classification errors in a “0/1” manner. For example, under standard metrics, predicting an “SUV” as a “car” will be penalized as much as erroneously labeling it as a “person”. This significantly deviates from the nuanced understanding humans apply when categorizing objects. To better align with humans, we adopt the recently proposed “open” metrics [58] for our upgraded benchmarks. In particular, for a groundtruth mask (pixel) $g_{i}$ with its label $c_{i}$ and the predicted mask (pixel) $d_{j}$ with its predicted label $c_{j}$, we compute soft $TP_{c_{i}}=S_{c_{i},c_{j}}$, soft $FP_{c_{j}}=1-S_{c_{i},c_{j}}$, and soft $FN_{c_{i}}=1-S_{c_{i},c_{j}}$, where $S_{c_{i},c_{j}}$ is the semantic similarity between $c_{i}$ and $c_{j}$. These soft TP, FP, and FN are then used to compute metrics PQ, AP, and mIoU in a standard way. With the “open” metrics, the evaluation of open-vocabulary segmentation would be more aligned with human perception, where closely related names incur lower penalties compared to those that are significantly different.
ADE20K [57] Cityscapes [7]
Benchmark Test Name Metric #classes PQ AP mIoU #classes PQ AP mIoU
Original Original Standard 150 24.30 16.79 32.14 19 39.42 22.76 53.08
Open [58] 150 27.04 16.52 36.59 19 42.57 26.54 62.69
OpenSeg Standard 150 25.35 17.30 33.06 19 43.68 26.93 56.15
Open [58] 150 29.03 17.52 41.03 19 45.92 27.09 60.21
Upgraded RENOVATE Standard 578 7.07 5.42 5.83 108 8.62 8.63 7.84
Open [58] 578 25.37 12.48 16.73 108 27.17 24.92 32.25
Table 2: Open-vocabulary segmentation on upgraded benchmarks. Our upgraded benchmarks with RENOVATE names contain more fine-grained annotations and are more challenging. Here we use pre-trained FC-CLIP model for evaluation.
In the original benchmarks, the evaluation metrics are computed on the original class division (150 classes for ADE20K and 19 classes for Cityscapes). In our upgraded benchmarks, the text prompt is the set of all the unique names in the datasets (578 classes for ADE20K and 108 classes for Cityscapes). When computing evaluation metrics, we also use this set of names as class division. Thus, the upgraded benchmarks contain much more classes than the original benchmarks. Since our names match better with the visual segments, the upgraded benchmarks are more realistically reflecting how human users may prompt the model for segmentation.
Results. In Tab. 2, we evaluate FC-CLIP model [52] on the upgraded benchmarks of ADE20K and Cityscapes and compare with the evaluation results on the original benchmarks. We also experiment with two other models, ODISE [48] and MasQCLIP [51], which lead to similar findings. See results in supplements.
We first observe that our benchmark is more challenging, as evidenced by considerably lower performance scores. This is because our benchmark requires the model to make classification on much more classes. This means predictions are closer to the decision boundaries and are thus easier to confuse between classes, leading to higher classification errors.
We also note that the discrepancy between standard and open metrics is much greater in our upgraded benchmarks. While standard benchmarks show minimal difference due to their class names being more distinct from each other, our benchmarks highlight a significant underestimation of the model’s open-vocabulary capability since the renovated names can be very similar. Open metrics address this by factoring in semantic similarity, providing a more realistic performance measure on our more fine-grained benchmarks.
Figure 6: Per-category IoU on the upgraded ADE20K with 578 classes. We highlight the top/bottom-5 renovated names and some selected names from the original “person” class. Our upgraded benchmarks allow us to conduct more fine-grained analysis on open-vocabulary model abilities and biases.
Finally, we demonstrate in Fig. 6 that our upgraded benchmarks enable a more detailed examination of models’ open-vocabulary capabilities and inherent biases. In particular, we find that the pre-trained model excels in identifying categories like “sky”, “girl”, and “sports car” with high mIoU scores, while struggling with “bollard”, “ceiling light”, and “tray”, reflecting the trend that large, well-defined, and common objects tend to be segmented more accurately than those that are small, deformable, and rare. Note that most of these names do not exist in the original benchmarks. In addition, our analysis reveals model biases, such as the disparity in mIoU scores between “man” and “woman”, “girl” and “boy”, suggesting an imbalance in the training dataset. Such fine-grained information is pivotal for improving models and informing dataset curation, demonstrating the substantial advantages of our upgraded benchmarks.
4.3 Training with renovated names
Setup. In this experiment, we investigate the benefit of using renovated names for training open-vocabulary models. Specifically, we train an FC-CLIP model on MS COCO with our renovated segment-matched names for the training set. We also train two other FC-CLIP models using the baseline names.
For a fair comparison of these models trained with different sets of names, during inference, we merge the three name sets to form a single class name set as the text prompt. For RENOVATE, the names for each class are simply the set of unique names for all the instances in that class. For evaluation, we group the name predictions based on the original class division. In this way, even when using more fine-grained prompts, we still compute the evaluation metrics for each original class. It’s important to note that our evaluation methodology aligns with those employed in previous studies [11, 48, 52] since the original and OpenSeg names also similarly accommodate multiple naming options for their classes. Since we compute metrics on the same class divisions as the original datasets, we only report standard metrics (PQ, AP, and mIoU) for comparing different models. We observe similar evaluation results when using open metrics. We detail them in the supplements.
Source dataset Target datasets
MS COCO ADE20K Cityscapes
Train names PQ AP mIoU PQ AP mIoU PQ AP mIoU
Original names 52.70 42.72 60.91 25.41 16.13 31.06 46.06 23.67 57.99
OpenSeg names 53.60 43.87 60.56 27.14 17.41 35.63 45.34 24.05 58.42
RENOVATE 56.62 45.50 65.48 27.98 17.94 37.50 46.10 27.99 61.25
Table 3: Training with renovated names. During inference, test names are merged from original, OpenSeg, and RENOVATE names for fair comparison. Our results demonstrate that RENOVATE names can help train stronger open-vocabulary models.
Results. Our results in Tab. 3 demonstrate that using RENOVATE names for training improves segmentation on both source dataset and target datasets. For example, the model trained with RENOVATE names obtains over 2% PQ on MS COCO, near 2% mIoU on ADE20K and around 3% AP and mIoU on Cityscapes. In addition, we notice that the mIoU metric typically gets improved more with our names than AP. This is consistent with our empirical findings that “stuff names” are typically more vague and thus have improvement potential than “thing names”. These results suggest that more precise and descriptive training names enhance the model generalization to both known and novel categories, potentially by better retaining the pre-trained vision-language alignment of the CLIP backbone during fine-tuning.
4.4 Automating name quality evaluation
Setup. We further leverage pre-trained open-vocabulary models as evaluators of names, automating the name quality evaluation process. Instead of involving humans, we prompt open-vocabulary segmentation models with different names. Names are then considered to be better if they help the model achieve better open-vocabulary segmentation. During inference, the text prompt consists of all the unique names from the corresponding name set. For evaluation, we adopt the same evaluation protocol as in Sec. 4.3 to compute metrics (PQ, AP, mIoU) on the same class divisions as the original datasets.
Panoptic Semantic
ADE-150 [57] Cityscapes [7] ADE-847 [57] PC-459 [27]
Model Test names PQ AP mIoU PQ AP mIoU mIoU mIoU
ODISE [48] Original names 21.88 13.94 29.16 39.72 27.73 49.60 10.75 16.97
OpenSeg names 21.63 13.95 28.88 43.26 28.45 54.53 11.15 17.81
RENOVATE 23.69 14.38 31.64 43.61 28.38 57.42 11.82 18.35
MasQCLIP [51] Original names 23.46 12.80 30.32 33.78 18.11 45.35 10.82 18.21
OpenSeg names 23.70 12.84 31.17 35.05 17.79 46.46 10.95 18.36
RENOVATE 25.00 12.93 32.30 35.63 18.19 50.07 11.39 18.85
FC-CLIP [52] Original names 24.30 16.79 32.14 39.42 22.76 53.08 13.33 18.32
OpenSeg names 25.35 17.30 33.06 43.68 26.93 56.15 13.78 19.22
RENOVATE 27.90 17.88 37.05 45.90 29.79 62.55 15.63 20.74
Table 4: Open-vocabulary segmentation evaluation. We use ODISE, MasQCLIP, and FC-CLIP pre-trained on COCO [22] and evaluate both panoptic and semantic segmentation performance. Our results show that RENOVATE class names are of high quality since they help boost the performance of open-vocabulary models.
Results. In Tab. 4, we experiment with three open-vocabulary segmentation models, i.e., MasQCLIP [51], ODISE [48], and FC-CLIP [52], all pre-trained on the MS COCO dataset. RENOVATE names consistently outperform the original and OpenSeg names across ADE20K, Cityscapes, and PASCAL Context datasets. Notably, FC-CLIP with RENOVATE names attains 15.6% and 16.4% relative PQ improvement from the original names on ADE20K and Cityscapes, respectively, setting new state-of-the-art performances on both benchmarks. This suggests that RENOVATE names align more accurately with the visual segments, enhancing segmentation performance. Note that some numbers in Tab. 4 are even better than when we train models with renovated names (Tab. 3). That is because the test names in Tab. 3 are of larger size with mixed name qualities from all three name sets, presenting a greater challenge. Despite this, when tested using the same names, as in Tab. 3, models trained with RENOVATE names are stronger than the others.
In addition, we observe again that mIoU is typically improved more with RENOVATE names than AP, further confirming our hypothesis that “stuff” class names have more improvement room. Moreover, we note that RENOVATE improves the performance of ODISE, which employs a Stable Diffusion [35] backbone instead of CLIP. This indicates the general applicability of RENOVATE names across different VLM backbones.
4.5 Ablation Analysis on the Renaming Pipeline
We further perform an ablation study using the same automated name quality evaluation as in Sec. 4.4. In Tab. 6, we first analyze the impact of context names for candidate generation. Our results indicate significant performance improvements when context information is provided, as opposed to the “None” approach, which relies only on original names. This highlights the effectiveness of leveraging contextual insights for generating superior candidate names. In comparison, CaSED outperforms both BLIP2 and RAM, achieving a greater boost in performance through its retrieval-based approach and the utilization of expansive external image-caption datasets, underlining the benefits of contextual enrichment in name generation.
Table 5: Ablation on the context name sources.
ADE20K Context name source PQ AP mIoU None 25.50 16.53 33.80 Caption nouns (BLIP2 [19]) 25.89 17.36 34.74 Image tags (RAM [55]) 26.95 17.55 35.54 Caption nouns (CaSED [6]) 27.90 17.88 37.05
ADE20K Attn bias Query Classifier PQ AP mIoU gt mask learnable text 21.09 14.91 26.96 gt mask text learnable 25.65 16.33 32.67 rand gt mask text learnable 26.21 17.50 35.62 rand gt mask text w/ neg learnable 27.90 17.88 37.05
Table 5: Ablation on the context name sources.
Table 6: Ablation on renaming model architecture.
In Tab. 6, we analyze the design choices of our renaming model. We start by establishing that our approach using text queries is superior to a standard text classifier, which only interacts with the transformer decoder’s output query embeddings and treats the mask prediction branch in a class-agnostic way. In comparison, our text queries interact intimately with spatial details from image features and exploit the mask prediction branch, leading to more precise name-segment matching. We also observe performance benefits from randomly substituting ground-truth masks with predicted ones in attention biases (“rand gt mask”), which suggests training with increased reliance on text queries refines the matching quality. Furthermore, augmenting our model with negative name appending (“text w/ neg”) enhances its discriminative capacity, further boosting performance and solidifying the renaming model’s overall effectiveness.
5 Discussion and Outlook
In this work, we highlight the problems of naming in existing open-vocabulary segmentation benchmarks and introduce RENOVATE, a straightforward yet versatile framework to renovate dataset names. Our efforts underscore the benefits of renaming, evidenced by the improvements in upgraded benchmarks and models trained with renovated names. However, while RENOVATE helps uncover biases in models and datasets, it could inadvertently propagate biases from the foundational models into the new names. To mitigate potential negative societal impacts, we advocate for verification of renovated names in critical applications, as we have exemplified with our upgraded benchmarks.
As the first attempt to propose a generalized, principle-based renaming framework, we acknowledge that our exploration of design choices remains incomplete. For instance, alternative language models like Llama [44] or Gemini [43] have yet to be investigated for candidate name generation, and our architectural experiments have so far been confined to using CLIP as the backbone, leaving other VLMs like SigCLIP [53] unexplored. In the future, we aim to further refine our renaming method, exploring a wider array of options, and ultimately scale it to automate the enhancement of publicly available large-scale datasets.
Supplementary Material
This supplementary material to the main paper “Renovating Names in Open-Vocabulary Segmentation Benchmarks” is structured as follows:
  • In Appendix 0.A, we provide a brief overview of works that address name problems in open-vocabulary classification.
  • In Appendix 0.B, we elaborate on the implementation details on the GPT-4 prompts for name generation, human verification process for benchmark upgrading, open metrics, human studies, and the renaming and pre-trained models.
  • In Appendix 0.C, we provide more experiment results on the upgraded benchmarks and on models trained with renovated names.
  • In Appendix 0.D, we present more examples of context, candidate, and renovated names. We also show more examples of name selection with the IoU scores.
  • In Appendix 0.E, we provide more qualitative examples demonstrating that RENOVATE can refine original names, correct wrong annotations, and uncover segments with shared semantic concepts across datasets.
Appendix 0.A A brief overview of name problems in open-vocabulary classification
While OpenSeg [11] is the only prior work for the name problems in open-vocabulary segmentation, there are more research works on improving name quality for open-vocabulary classification. For example, [10] improve class names by including their parents and children from the WordNet hierarchy [26].  [6] propose to replace the fixed vocabulary list with nouns extracted from best-matching captions.  [24] improve the descriptiveness of the vocabulary by adding class descriptions. However, since the VLMs used by these models are all trained on image-text datasets, it is non-trivial to extend them to dense prediction tasks like segmentation. For the renaming strategy, our context name design and LLM-based generation approach ensure the diversity of our names with hierarchical concepts as well as descriptive contexts, without the use of an external word hierarchy as [26] or class descriptions as [24]. Our renaming model further improves the name quality based on the visual segments and is more specifically designed for the segmentation tasks.
Appendix 0.B More implementation details
0.B.1 GPT-4 prompts for name generation
We use GPT-4 ChatCompletion API [1] for candidate name generation. For each class, we provide two messages. The first message, i.e., “system message”, describes the role of the system, where we describe our overall problem setup and goals. The second message, i.e., “user message”, describes the intention of the user, where we provide the original class names and context names. We show our exact prompt messages as follows:
System message: You are a helpful assistant aiding in renaming dataset classes. Each class has an inadequate original name and a set of context names derived from related captions (with their frequencies sorted and listed in brackets). These context names provide insights into the category’s essence. When renaming, you may: 1. Use synonyms or subcategories of the original class name (e.g., ‘grass’ can be renamed as ‘lawn, turf’). 2. Provide a short context to address polysemy (e.g., ‘fan’ can be renamed as ‘ceiling fan, floor fan’). Please generate new names for each class in lower case, listed in a row. Ensure the new names logically connect to the original class, using it as the head noun. Avoid arbitrary noun concatenations and nonsensical names. For instance, the class ‘sky’ should not yield names like ‘person under sky’. Ready to proceed with naming? Kindly provide the original class names and context names.
User message: Original name: {original class name}, context names (with frequencies) are {context names}. What are the new names? Only provide 5-10 names. And make sure you generate at least 3 reasonable synonyms or subcategories.
For the system message, we start the message with “You are a helpful assistant” following the examples in OpenAI Chat Completions API [1]. We then describe our problem setup, input format, and instructions with examples on name generation. For the user message, {original class name} and {context names} require inputs from each category. For our ablation experiments without context names (“None” in Tab. 6 in the main paper), we remove the descriptions about context names in the system message and do not require context names in the user message.
0.B.2 More details on the human verification process
We upgrade the benchmarks on ADE20K and Cityscapes with our renovated names. To ensure the high quality of names, we verify the names on all the segments in their validation set. As shown in Fig. 0(a), for each segment, we provide the model-selected name along with the other top 3 names ranked by the model, we also provide the rest of the candidate names in the “others” option. Therefore, the human verifier can quickly continue to the next segment if the selected name itself or another name in top 3 names matches the most with the segment. If none of the top 3 names fits well, humans need to choose from the other candidate names in “others”. In the latter case, it would be similar to the scenario where RENOVATE suggestions are not provided and humans have to choose from the whole un-ranked candidate name list. In Fig. 0(b), we compare the name verification speed with and without RENOVATE suggestions on 200 segments. As we can see, RENOVATE suggestions significantly speed up our verification process. We note that the human verifiers are provided only with candidate names from the same original class, which is typically the original ground-truth class. In the rare cases where our renaming model prefers names from classes other than the original groundtruth class, we first verify the class choices, then provide the human verifiers with the name suggestions from the verified class.
We further note that for all the segments verified, 68% of them are matched to the default, i.e., top 1, model-selected name, and 92% of them are matched to the top 3 RENOVATE suggestions, reflecting the high quality of RENOVATE names. In Fig. 0.B.2, we show typical cases when humans and models disagree, including segments that are too small to recognize and segments whose names cannot be inferred correctly due to insufficient visual cues. We note that such segments only consist a minority of all the segments, with inherent difficulty to name precisely.
(a) User interface for name verification.
(b) Comparison of verification speed.
Figure 0.B.1: Human verification for upgrading benchmarks. We ask 5 human annotators to verify the names of segments in the validation sets of ADE20K and Cityscapes. As shown in (a), humans are given the top 3 suggestions from the model and are asked to verify the selected name or choose a more matching name from either the other top 3 names or the rest of the candidate names. (b) shows that RENOVATE name suggestions significantly speed up the human verification/annotation process.
(a) Segments are too small.
(b) Segments lack sufficient visual cues.
Figure 0.B.2: Typical cases when humans and models disagree during human verification. We show two typical cases when humans may choose different names from our model-selected names. (a) shows when segments are too small to recognize, it can be difficult to decide which name is correct. This may indicate a limitation of our renaming method on extremely small objects. (b) shows when there is a lack of sufficient visual cues to infer the names of the segments, the choices are ambiguous. For example, both “river” and “lake” are reasonable choices for the segment in the top right image without more information of the scene. Note (a) and (b) consist only a minority of all the segments.
0.B.3 More details on the open metrics in upgraded benchmarks
To compute the open metrics [58], we need a similarity function on names. Previous research [58] compared three types of similarity measures: WordNet methods [18, 46, 21], text embedding models [25, 31, 15], and pre-trained language models [32, 8, 33, 34]. They demonstrated that both WordNet methods and text embedding models compute reasonable similarity measures while pre-trained language models tend to overestimate similarity between names. However, since the names in our work are in free form and may not map to any synonym sets in WordNet [26], we adopt one of the text embedding models, FastText [15], which is shown to compute reasonable similarity measures [58].
0.B.4 Other implementation details
More details on the human study. The human study (Sec. 4.1) is conducted on ADE20K and Cityscapes datasets, which consists of 30249 and 14670 segments in the validation set, respectively. We randomly select 10% of the validation set for human study. The average time for human experts to finish the preference study is about 2 hours.
More details on the renaming model. For the backbone of the renaming model, we use ConvNeXt-Large CLIP [32, 13] from OpenCLIP [13] pretrained on LAION-2B dataset [38]. Following Mask2Former [4], our transformer decoder consists of nine blocks and we use AdamW [23] optimizer with a learning rate of $10^{-4}$ and a multi-step decay schedule with weight decay of 0.05.
More details on the pretrained models. ODISE [48] provides two pretrained models, ODISE (label) and ODISE (caption), trained on category names and caption nouns respectively. We use ODISE (label) since they demonstrated stronger performance than the other model in the original paper. For ODISE and FC-CLIP, we set their geometric mean parameters $\alpha=\beta=0.5$ so we do not differentiate the treatment of test names based on their name overlapping with the training names. In this way, we can compare different test names in a fair way. By default, we use FC-CLIP for all of our ablation studies. For the test prompt template, we follow prior work [12, 51, 52] to use the ViLD template [12], which consists of 63 prompt templates such as “There is {article} {category} in the scene.” These templates are ensembled by averaging the text embeddings.
Appendix 0.C More experiment results
ADE20K [57] Cityscapes [7]
Benchmark Test Name Metric #classes PQ AP mIoU #classes PQ AP mIoU
Original Original Standard 150 21.88 13.94 29.16 19 39.72 27.73 49.60
Open [58] 150 26.31 14.49 34.39 19 44.09 27.91 58.20
OpenSeg Standard 150 21.63 13.95 28.88 19 43.26 28.45 54.53
Open [58] 150 28.02 15.32 38.86 19 46.66 28.88 58.61
Upgraded RENOVATE Standard 578 3.15 1.21 5.64 108 5.93 4.54 6.25
Open [58] 578 10.35 4.03 22.13 108 18.27 15.15 20.20
Table 0.C.1: Open-vocabulary segmentation on upgraded benchmarks using ODISE [48]. Similar to Tab. 2, the upgraded benchmarks are considerably more challenging and the gap between standard and open metrics is significantly larger on the upgraded benchmark, highlighting the need of open metrics for more human-aligned evaluation when names are more similar.
ADE20K [57] Cityscapes [7]
Benchmark Test Name Metric #classes PQ AP mIoU #classes PQ AP mIoU
Original Original Standard 150 23.46 12.80 30.32 19 33.78 18.11 45.35
Open [58] 150 26.53 13.01 36.33 19 37.94 18.35 55.75
OpenSeg Standard 150 23.70 12.84 31.17 19 35.05 17.79 46.46
Open [58] 150 28.04 13.89 39.40 19 38.83 18.62 52.11
Upgraded RENOVATE Standard 578 6.36 3.78 3.90 108 6.12 4.58 5.53
Open [58] 578 24.27 10.78 18.13 108 16.50 15.42 18.89
Table 0.C.2: Open-vocabulary segmentation on upgraded benchmarks using MasQCLIP [51].
More results on upgraded benchmarks. In addition to FC-CLIP model [52] (Sec. 4.2), we further evaluate two other models on the upgraded benchmarks of ADE20K and Cityscapes and compare with the evaluation results on the original benchmarks. The results in Tab. 0.C.1 and Tab. 0.C.2 are consistent with our findings in Sec. 4.2. Specifically, since names in our upgraded benchmarks are more fine-grained and more similar, our benchmarks are more challenging and the discrepancy between standard and open metrics is much greater.
More results on training with renovated names. We further evaluate the models trained with different sets of names using open metrics in Tab. 0.C.3. The results are consistent with Tab. 3 and demonstrate that training with RENOVATE names improves the model’s open-vocabulary generalization ability.
Source dataset Target datasets
MS COCO ADE20K Cityscapes
Train names PQ AP mIoU PQ AP mIoU PQ AP mIoU
Original names 53.55 42.69 64.81 28.24 16.70 38.34 47.57 26.94 63.13
OpenSeg names 55.05 43.47 64.52 30.55 17.84 42.18 46.93 27.51 62.48
RENOVATE 57.34 46.44 66.47 30.89 18.10 44.24 47.90 30.14 65.35
Table 0.C.3: Evaluating models trained with different names using open metrics. Except for the metric, the evaluation process is the same as in Sec. 4.3.
Original name Context names Candidate names Renovated names mIoU gain earth, ground building, tree, stand, water, person, floor, sky, road, snow, car ground cover,earth floor,ground cover, ground,earth surface,snow ground, land,natural ground,outdoor ground,terrain, earth,dirt ground,terrestrial terrain ground cover, terrestrial terrain, natural ground, snow ground, dirt ground, ground, grass +11.50 field lush, field, sky, green, grass, tree, road, hillside, grassy, rural rural field, roadside field, green field, crop field, sports field, grassland, grassy hillside, rural field, roadside field, grassland, crop field, sports field +15.02 rock, stone stone, water, tree, lush, sea, sky, shoreline, coast, creek, stand rock,stone,standing stone,beach pebble, shoreline rock,national park boulders, river stone,stones,lush stone,rocks river stone, stones, beach pebble rock, national park boulders, shoreline rock +8.95 fountain city, fountain, building, garden, square, park, water, hotel, place, design garden fountains,fountain,urban fountains, square fountains,park fountains, water feature designs,hotel water features hotel water features, urban fountains,park fountains +31.58 counter room, area, lobby, reception, shop, store, hotel, design, office, interior checkout counter,reception desk, office counter area,countertop,hotel lobby counter, cashier counter,counter,retail store counter, service desk,storefront counter,information kiosk information kiosk, reception desk, cashier counter, counter, office counter area, checkout counter +20.72 stairway, staircase building, stair, floor, room, ceiling, lead to, pillar, balustrade, stairwell, rail balustrade staircase,ceiling stairwell,stairway, building stairs,staircase,pillar-supported stairs, room stairway,railing stairway,floor staircase pillar-supported stairs,railing stairway, room stairway, floor staircase +14.96 bar bar, restaurant, room, area, hotel, interior, pub, table, house, counter bar table,bar area,restaurant bar,hotel bar, home bar,bar counter,bar interior,bar,pub bar bar area, bar table bar counter +18.58 canopy room, bed, bedroom, house, idea, design, palace, garden, home, royal home canopy,canopy,house canopy, garden canopy,royal palace canopy,patio canopy, bedroom canopy,room canopy bedroom canopy, patio canopy, garden canopy +9.88 grandstand, covered stand stadium, game, field, arena, world, park, football, school, sport, high high-capacity grandstand, park bleachers, covered game stand, world-class viewing area, spectator stand, football field seating,grandstand, sports arena stand,stadium grandstand, school spectator seats sports venue grandstand, sports arena stands, park bleachers, spectator stand football field seating +29.50 swimming pool, swimming bath, natatorium pool, swimming, house, hotel, villa, resort, beach, indoor, luxury, apartment luxury resort pool,swimming bath,villa pool, apartment complex pool,hotel swimming pool, home swimming pool,swimming pool, beachfront swimming pool,indoor swimming bath hotel swimming pool, beachfront swimming pool, villa pool, indoor swimming pool +30.12
Table 0.C.4: Examples of context names, candidate names, and RENOVATE names for classes from ADE20K.
Figure 0.C.1: More examples on name selection based on IoU scores.
Appendix 0.D Examples of context, candidate, renovated names, and name selection
In Tab. 0.C.4, we show examples of original names, context names, candidate names, RENOVATE names, and their mIoU gain in the open-vocabulary segmentation evaluation as in Sec. 4.4. For RENOVATE, the names for each class are simply the set of unique names for all the instances in that class after renaming. We can see that the context names are important for understanding the vague class names such as “earth” and “field” and are helpful for providing more contexts to names with multiple potential meanings like “counter” and “bar”. We can also see that RENOVATE names provide more comprehensive descriptions of the corresponding classes. Using these names to prompt open-vocabulary models leads to significant performance gains, suggesting that our renovated names match better with the visual segments.
In Fig. 0.C.1, we further show more visual examples of the name selection process (Sec. 3.3). Together with Fig. 3 in the main paper, these examples demonstrate that our renaming model provides a meaningful ranking of the names to match with each segment.
Appendix 0.E More qualitative analysis
0.E.1 RENOVATE can refine original names
In Fig. 0.E.1 and Fig. 0.E.2, we present more examples of RENOVATE names along with the original names of segments from ADE20K and MS COCO. We show that RENOVATE refines the original names by being more precise and closer to human natural language.
Figure 0.E.1: More examples of renovated names on segments from the validation set of ADE20K. For each segment, we show the original name below the image and the renovated name in the text box.
Figure 0.E.2: More examples of renovated names on segments from the validation set of MS COCO. For each segment, we show the original name below the image and the renovated name in the text box.
0.E.2 Other potential use cases of RENOVATE
RENOVATE can correct annotation errors. From Fig. 0.E.3, we can see that our renaming model discovers wrong annotations and finds more accurate names through RENOVATE. It’s important to highlight that these renovated names do not belong to the original ground-truth class; instead, our model ranks them over names from the original class. This demonstrates the renaming model’s utility in detecting and correcting inaccuracies in annotations by effectively matching names with visual content.
Figure 0.E.3: RENOVATE can find wrong annotations and suggest corrections.
RENOVATE uncovers segments with shared semantic concepts across datasets. Since RENOVATE aims to describe visual segments using more human-aligned language, it effectively standardizes the naming conventions for identical concepts in different datasets. In Fig. 0.E.4, we demonstrate this with examples from MS COCO and ADE20K. While renaming each dataset separately, we observe that RENOVATE inadvertently reveals the shared name spaces by identifying more appropriate names in natural language to describe the visual contents. This property can be potentially used for analyzing the differences between datasets, merging datasets to construct new benchmarks, and various other data curation processes.
Figure 0.E.4: RENOVATE uncovers segments with shared semantic concepts across datasets.